Find limit using L' Hopitals rule:


(a) lim x "\to" 0 (cos2x)^(1/x^2)

1
Expert's answer
2021-09-28T09:18:04-0400

(a) "lim_{x\rightarrow 0} (cos2x)^{\frac{1}{x^2}}"

"\because u=e^{ln(u)}"


"\implies lim_{x\rightarrow0}(cos2x)^{1/x^2}=lim_{x\rightarrow0}e^{ln(cos2x)^{1/x^2}}=lim_{x\rightarrow0}e^{\frac{ln(cos2x)}{x^2}}"


"\implies" Since we have an intermediate of the type "\dfrac{0}{0}" , So we can use L'Hopital's Rule -


"\implies e^{lim_{x\rightarrow0}\frac{cos2x}{x^2}}=e^{lim_{x\rightarrow0}\frac{\frac{d}{dx}(ln(cos2x))}{\frac{d}{dx}(x^2)}} = e^{lim_{x\rightarrow0}\frac{(-sin2x)}{xcos2x}} = e^{lim_{x\rightarrow0}(\frac{-tan2x}{x})}"


"\implies e^{-lim_{x\rightarrow0}(\frac{tan2x}{x})}"


Again, Since we have an intermediate of the type "\dfrac{0}{0}" , So again we can use l'Hopital's Rule -

"\implies e^{-lim_{x\rightarrow0}(\frac{tan2x}{x})}" "= e^{-lim_{x\rightarrow 0}(\frac{d/dx(tan2x)}{d/dx(x)})}=e^{-lim_{x\rightarrow 0}(2sec^22x)}"


"\\\implies" "e^{-lim_{x\rightarrow 0}(2sec^22x)} = e^{-2\times 1}=e^{-2}"


Hence, "lim_{x\rightarrow 0 }(cos2x)^{\frac{1}{x^2}}=\dfrac{1}{e^2}"